# Small solutions of ternary quadratic congruences with averaging over the moduli

### v3

@@ 1. Introduction and main results
In our recent paper [1], we proved the following almost-all result on small solutions of ternary quadratic congruences (see [1, first part of Theorem 1 in a slightly modified form]).
@@
This was a direct consequence of the following asymptotic result on the number of solutions to (1) of bounded heighth (see [1, second part of Theorem 1]).
@@
$\mathcal{B}(Q,\alpha_{2}):=\{(s,t)\in\mathbb{N}^{2}:Q<s\leq 2Q,\ \mbox{\rm gcd}(2\alpha_{2},s)=1,\ 1\leq t\leq s,\ \mbox{\rm gcd}(s,t)=1\}.$
@@
The novelty in this paper, as compared to [1], is to use an eighth moment bound for short character sums to moduli in a dyadic interval. We will be brief at places where the arguments from [1] repeat. Throughout the sequel, $\varepsilon$ will be an arbitrarily small positive number which, depending on the context, may change from a line to the next. For $n\in\mathbb{N}$, we denote by $\tau(n)$ the number of divisors of $n$ and recall the well-known bound $\tau(n)\ll_{\varepsilon}n^{\varepsilon}$.
@@ 2. Preliminaries
This is [4, Theorem 12.6] if $\chi$ is primitive. If $\chi$ is non-primitive and induced by a primitive character $\chi_{1}$ of conductor $q_{1}>1$ dividing $q$, then
@@
Suppose $f:(0,\infty)\to\mathbb{C}$ is a continuous function of bounded variation and $a<b$ are real numbers such that the integral defining its Mellin transform $\hat{f}(s)$ converges absolutely for every $s$ in the strip $\{z\in\mathbb{C}:a<\Re z<b\}$. Then for every $x>0$ and $c\in(a,b)$, the inversion formula
@@
holds, where the integral is taken along the vertical line $\Re(s)=c$.
@@
Let us first assume that $\chi$ is a primitive character of conductor $q>1$. Then for $0\leq\sigma\leq 1$, the claimed bound (10) follows from [4, Exercise 3 on page 100], applied to Dirichlet $L$-functions. For $\sigma\geq 1$, this bound is standard and can be achieved via partial summation. If $\chi$ is non-primitive and induced by a primitive character $\chi_{1}$ of conductor $q_{1}>1$ dividing $q$, then
@@
if $\sigma=\Re s\geq 0$, and hence, the claimed bound (10) remains valid in this case. ∎
@@
This follows from [3, Proposition 3.2 with $c=0$]. ∎
@@
Tacitly, we will use the well-known bounds
$\frac{q}{\varphi(q)}\ll\log\log(10q)\quad\mbox{and}\quad\sum\limits_{Q<q\leq 2Q}\frac{1}{\varphi(q)^{k}}\ll_{k}Q^{1-k}\quad\mbox{for }k\in\mathbb{N}$
@@ 3. Initial steps
Proceeding along the lines of [1, section 4], we obtain the asymptotic formula
@@
As in [1, section 3], we use Dirichlet characters to detect the congruence condition $x_{1}^{2}+\alpha_{2}x_{2}^{2}+\alpha_{3}x_{3}^{2}\equiv 0\bmod{q}$, arriving at
@@
In [1, sections 6 and 7], we demonstrated that (with 1 in place of the summand $\Phi(x_{3})$ and $\varepsilon$ in place of $3\varepsilon$)
@@ 4. Elementary estimates
$F:=\sum_{\begin{subarray}{c}Q<q\leq 2Q\\ \text{gcd}(2\alpha_{2},q)=1\end{subarray}}\frac{1}{\varphi(q)}\sum\limits_{\begin{subarray}{c}\chi\bmod{q}\\ \chi^{2}\neq\chi_{0}\end{subarray}}\bigg|\sum_{x_{3}\in\mathbb{Z}}\Phi(x_{3})\chi^{2}(x_{3})\bigg|^{8}.$
The following two lemmas provide bounds for $E_{1}$ and $E_{2}$.
@@
This follows from [1, equations (19) and (20)]. ∎
@@
(20) $G_{1}:=\sum\limits_{\begin{subarray}{c}|x_{1}|,|x_{2}|,|y_{1}|,|y_{2}|\leq N\\ |u_{1}|,|u_{2}|,|v_{1}|,|v_{2}|\leq N\\ (x_{1}^{2}+\alpha_{2}x_{2}^{2})(y_{1}^{2}+\alpha_{2}y_{2}^{2})\neq(u_{1}^{2}+\alpha_{2}u_{2}^{2})(v_{1}^{2}+\alpha_{2}v_{2}^{2})\end{subarray}}\sum\limits_{q|\left((x_{1}^{2}+\alpha_{2}x_{2}^{2})(y_{1}^{2}+\alpha_{2}y_{2}^{2})-(u_{1}^{2}+\alpha_{2}u_{2}^{2})(v_{1}^{2}+\alpha_{2}v_{2}^{2})\right)}1.$
Recall that $\alpha_{2}$ is assumed to be a positive integer. Applying the divisor bound $\tau(n)\ll_{\varepsilon}n^{\varepsilon}$ gives
@@ 5. Estimation of $F$
The integrand is analytic in the half plane $\Re s>0$, and hence, we may use Cauchy’s integral theorem to write
@@
(23) $\hat{\Phi}(\sigma+it)\ll_{j}|t|^{-j}N^{\sigma-1+j}\int\limits_{\mathbb{R}}|\Psi^{(j)}(x)|dx\ll\Delta N^{\sigma-1}\cdot\left(\frac{N}{\Delta|t|}\right)^{j},$
@@
(25) $c:=1+\frac{1}{\log 2N}\quad\mbox{and}\quad T:=\frac{Q^{\varepsilon}N}{\Delta}.$
@@ 6. Proofs of Theorems 4 and 5
Remark 2: If we take $r=4$ in place of $r=3$ in our application of Proposition 6 above, then we obtain the slightly better exponent $159/368$ in place of $45/104$. However, in this case, the said Proposition 6 is applicable only if $q$ is cubefree.

### v4

@@ 1. Introduction and main results
In this paper, we continue our study of small solutions of diagonal ternary quadratic congruences
(1) $\alpha_{1}x_{1}^{2}+\alpha_{2}x_{2}^{2}+\alpha_{3}x^{3}\equiv 0\bmod{q},$
begun in [1] and [2].
The study of quadratic congruences of the form $Q(x_{1},...,x_{n})\equiv 0\bmod{q}$, with $Q$ an integral quadratic form, has attracted a lot of attention. In particular, if $q$ is odd and squarefree, Heath-Brown [6, Theorem 2] proved that for any integral ternary quadratic form $Q(x_{1},x_{2},x_{3})$ with determinant coprime to $q$, there exists a solution $(x_{1},x_{2},x_{3})\in\mathbb{Z}^{3}\setminus\{(0,0,0)\}$ of height $\max\{x_{1},x_{2},x_{3}\}\ll q^{5/8+\varepsilon}$ to the congruence
(2) $Q(x_{1},x_{2},x_{3})\equiv 0\bmod{q}$
(see [3] for a corresponding result regarding the case of prime power moduli). It may be expected that the exponent $5/8$ in Heath-Brown’s result can be replaced by $1/2$. In general it is not possible to lower this exponent further as the example of the congruence $x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\equiv 0\bmod{q}$ shows. However, we proved in [2] that, in a sense, almost all congruences of the form in (1) modulo a fixed $q$ have a solution whose height is much smaller than $q^{1/2}$. Precisely, we established the following result (see [2, first part of Theorem 1 in a slightly modified form]).
@@
This was a direct consequence of the following asymptotic result on the number of solutions to (3) of bounded heighth (see [2, second part of Theorem 1]).
@@
$\mathcal{B}(Q,\alpha_{2}):=\{(q,t)\in\mathbb{N}^{2}:Q<q\leq 2Q,\ \mbox{\rm gcd}(2\alpha_{2},q)=1,\ 1\leq t\leq q,\ \mbox{\rm gcd}(q,t)=1\}.$
@@
The novelty in this paper, as compared to [2], is to use an eighth moment bound for short character sums to moduli in a dyadic interval. We will be brief at places where the arguments from [2] repeat. Throughout the sequel, $\varepsilon$ will be an arbitrarily small positive number which, depending on the context, may change from a line to the next. For $n\in\mathbb{N}$, we denote by $\tau(n)$ the number of divisors of $n$. Recall the well-known bound $\tau(n)\ll_{\varepsilon}n^{\varepsilon}$ for every $\varepsilon>0$.
@@ 2. Preliminaries
This is [8, Theorem 12.6] if $\chi$ is primitive. If $\chi$ is non-primitive and induced by a primitive character $\chi_{1}$ of conductor $q_{1}>1$ dividing $q$, then
@@
Suppose $f:(0,\infty)\to\mathbb{C}$ is a continuous function of bounded variation and $a<b$ are real numbers such that the integral defining its Mellin transform $\hat{f}(s)$ converges absolutely for every $s$ in the strip $\{z\in\mathbb{C}:a<\mbox{Re}(z)<b\}$. Then for every $x>0$ and $c\in(a,b)$, the inversion formula
@@
holds, where the integral is taken along the vertical line $\mbox{Re}(s)=c$.
@@
Let us first assume that $\chi$ is a primitive character of conductor $q>1$. Then for $0\leq\sigma\leq 1$, the claimed bound (12) follows from [8, Exercise 3 on page 100], applied to Dirichlet $L$-functions. For $\sigma\geq 1$, this bound is standard and can be achieved via partial summation. If $\chi$ is non-primitive and induced by a primitive character $\chi_{1}$ of conductor $q_{1}>1$ dividing $q$, then
@@
if $\sigma=\mbox{Re}(s)\geq 0$, and hence, the claimed bound (12) remains valid in this case. ∎
@@
This follows from [7, Proposition 3.2 with $c=0$]. ∎
@@
Tacitly, we will use the well-known bound (see [5, Theorem 328])
$\frac{q}{\varphi(q)}\ll\log\log(10q)$
@@ 3. Simplification of the problem
Proceeding along the lines of [2, section 4], we obtain the asymptotic formula
@@
As in [2, section 3], we use Dirichlet characters to detect the congruence condition $x_{1}^{2}+\alpha_{2}x_{2}^{2}+\alpha_{3}x_{3}^{2}\equiv 0\bmod{q}$, arriving at
@@
In [2, sections 6 and 7], we demonstrated that (with 1 in place of the summand $\Phi(x_{3})$ and $\varepsilon$ in place of $3\varepsilon$)
@@ 4. Elementary estimates
$F:=\sum_{\begin{subarray}{c}Q<q\leq 2Q\\ \text{gcd}(2\alpha_{2},q)=1\end{subarray}}\frac{1}{\varphi(q)}\sum\limits_{\begin{subarray}{c}\chi\bmod{q}\\ \chi^{2}\neq\chi_{0}\end{subarray}}\bigg|\sum_{x_{3}\in\mathbb{Z}}\Phi(x_{3})\chi^{2}(x_{3})\bigg|^{8},$
where we have replaced $\overline{\chi}$ by $\chi$. The following two lemmas provide bounds for $E_{1}$ and $E_{2}$.
@@
This follows from [2, equations (19) and (20)]. ∎
@@
(22) $G_{1}:=\sum\limits_{\begin{subarray}{c}|x_{1}|,|x_{2}|,|y_{1}|,|y_{2}|\leq N\\ |u_{1}|,|u_{2}|,|v_{1}|,|v_{2}|\leq N\\ (x_{1}^{2}+\alpha_{2}x_{2}^{2})(y_{1}^{2}+\alpha_{2}y_{2}^{2})\neq(u_{1}^{2}+\alpha_{2}u_{2}^{2})(v_{1}^{2}+\alpha_{2}v_{2}^{2})\end{subarray}}\sum\limits_{\begin{subarray}{c}q\in\mathbb{N}\\ q|\left((x_{1}^{2}+\alpha_{2}x_{2}^{2})(y_{1}^{2}+\alpha_{2}y_{2}^{2})-(u_{1}^{2}+\alpha_{2}u_{2}^{2})(v_{1}^{2}+\alpha_{2}v_{2}^{2})\right)\end{subarray}}1,$
where we have interchanged summations and then dropped the condition $Q<q\leq 2Q$ in $G_{1}$. Recall that $\alpha_{2}$ is assumed to be a positive integer. Applying the divisor bound $\tau(n)\ll_{\varepsilon}n^{\varepsilon}$ gives
@@ 5. Estimation of the main sum
The integrand is analytic in the half plane $\mbox{Re}(s)>0$, and hence, we may use Cauchy’s integral theorem to write
@@
(25) $\hat{\Phi}(\sigma+it)\ll_{j}|t|^{-j}N^{\sigma-1+j}\int\limits_{\mathbb{R}}|\Phi^{(j)}(x)|dx\ll\Delta N^{\sigma-1}\cdot\left(\frac{N}{\Delta|t|}\right)^{j},$
@@
(27) $c:=1+\frac{1}{\log(2N)}\quad\mbox{and}\quad T:=\frac{Q^{\varepsilon}N}{\Delta}.$
@@ 6. Proofs of Theorems 4 and 5
Remark 2: If we take $r=4$ in place of $r=3$ in our application of Proposition 6 above, then we obtain the slightly better exponent $159/368$ in place of $45/104$. However, in this case, the said Proposition 6 is applicable only if $q$ is cubefree. Taking larger values $r>4$ in Proposition 6 does not improve the result.